OpenStudy (anonymous):

Solve using substitution: x^2+3y^2=172 2x-y=3

OpenStudy (anonymous):

Get y by itself in second equation

OpenStudy (anonymous):

y = 2x-3. use that in the first equation. you will get a quadratic in x. solve for x and use those values of x to find y.

OpenStudy (anonymous):

y=-3-2x substitute y in the first equation

OpenStudy (anonymous):

x^2+3(-3+2x)^2=172 2(11.15)-y=3 x^2+3(9+4x^2)=172 22.3-y=3 x^2+27+12x^2=172 -y=3-22.3 13x^2=172-27 -y/-1=-19.3/-1 13x^2=145 y=19.3 13x^2/13=145/13 x^2=11.15384615\[\sqrt{?}\] x=11.15